A 20 mM solution of ethanol is diluted 1 in 8. The diluted ethanol solution was used for an alcoholdehydrogenase assay that had a total volume of 1.5 mL. 20 μL of alcoholdehydrogenase was used for the assay. The reaction was followed spectrophotometrically and a progress curve plotted from the spectrophotometric data. Using the progress curve, the initial velocity was calculated as 0.30 ΔA/min.


The assay was repeated as above, with a different volume of alcoholdehydrogenase . The initial velocity was 0.01 ΔA/s. What volume (in μL) of alcoholdehydrogenase was used for the assay?


1
Expert's answer
2021-09-17T02:08:53-0400

Volume of alcoholdehydrogenase in first experiment, V1 = 20 μL

Initial velocity in first experiment, U1 = 0.30 ΔA/s

Initial velocity in second experiment, U2 = 0.01 ΔA/min = 0.01 ΔA/min / 60 (Dividing by 60 to convert min to sec)

Initial velocity in second experiment, U2= 0.00016 ΔA/s

Volume of alcoholdehydrogenase in second experiment, V2 = (V1 * U1) / U2 = (20μL * 0.30 ΔA/s) / 0.00016 ΔA/s = 37500μL

Volume of alcoholdehydrogenase used for the assay = 37500 μL